# All-DNA associative polymers

### v1

@@ I Introduction
In recent years, associative polymers have been extensively employed as model systems to understand the behavior of intrinsically disordered proteins (IDPs) and other biopolymers, which often rely on multivalent, weak interactions to drive phase separation and compartmentalization without membranes [11, 12]. To capture these complex behaviors, associative polymers are typically modeled using coarse-grained representations, where the backbone is treated as a polymer chain and the stickers as discrete sites capable of specific pairwise interactions [1, 2]. In the case of simple models, where chains are flexible, spacers interact through excluded volume only, and each sticker can be involved in no more than one bond, there is theoretical [3], numerical [13, 14, 15], and experimental [7, 16] evidence that no thermodynamic instability is encountered upon increasing concentration. Indeed, in the dilute regime most of the bonds involve stickers belonging to the same chain (intra-molecular bonds), which therefore resembles a single-chain nanoparticle [17]. However, as concentration increases, chains start to interact more strongly with each other, forming inter-molecular connections that are entropically favoured [18, 19], and therefore progressively replace the intra-molecular bonds, thereby giving raise to a network-like structure without any phase separation occurring. However, it has been recently shown that in simple bead-spring models the different entropic contributions that control the overall bonding process can be manipulated by introducing multiple sticker types, with the constraint that only like stickers can bind to each other. Interestingly, as soon as there are at least two sticker types alternating along the backbone, an entropy-driven phase separation appears for a wide range of system parameters [20, 21]. The main contribution dictating the phase behaviour is due to the much larger entropic cost of closing an intra-molecular bond compared to an inter-molecular bond, effectively enhancing the number of inter-molecular contacts, and therefore the overall tendency to phase separate.
@@ II Methods
We simulate two one-component model systems made of single strand DNA in solution that we will call $(AA)_{6}$ and $(AB)_{6}$. Each DNA strand is composed by $N_{\rm nucl}=158$ nucleotides, with a backbone of thymine bases (the spacers) in which are inserted $M=12$ self-complementary regions (the stickers) as shown in Fig. 1, which also reports the full sequences of the strands used throughout this work. In the $(AA)_{6}$ systems the stickers have all the same sequence (TCGA), while in the $(AB)_{6}$ systems the stickers have alternating TGCA and GTAC sequences. The two sequences are self-complementary, so that only stickers of the same type can bind to each other.
@@
[…]
@@
During all simulations we keep track of pairs of nucleotides sharing hydrogen bonds (HBs), where two nucleotides are considered bonded if their hydrogen bond interaction energy is lower than $-0.1$ in oxDNA internal units, classifying the bond as intra- or inter-molecular if the two nucleotides belong to the same strand or to different strands, respectively. We then consider two stickers as bonded if their nucleotides are involved in at least two bonds. We thus obtain the total number of stickers involved in intra- and inter-molecular bonds, $N_{\rm intra}$ and $N_{\rm inter}$, and define the fraction of stickers involved in intra- and inter-molecular as $f_{\rm intra}=N_{\rm intra}/N_{\rm max}$ and $f_{\rm inter}=N_{\rm inter}/N_{\rm max}$, where $N_{\rm max}=MN_{s}$ is the total number of stickers. We also define the fraction of bonded stickers as $f_{\rm tot}=(N_{\rm intra}+N_{\rm inter})/N_{\rm max}$.
@@ III Results
Figure 2: The fraction of hydrogen bonds (HBs) between nucleotides belonging to stickers involved in total, intra- or inter-molecular bonds for the (left) $(AA)_{6}$ and (right) $(AB)_{6}$ systems. Each row represents a different density which takes the values, from top to bottom, of $\rho\sigma^{3}=0.57$, $\rho\sigma^{3}=0.31$, and $\rho\sigma^{3}=0.16$, respectively.
@@
Figure 3: (3(a)) Graphic representation of the system in a direct-coexistence simulation at the beginning of the calculation. (3(b)) Density profiles at the start of the simulation (dashed black curve) and at the end of the simulation, when the system has reached equilibrium (solid orange), for the $(AA)_{6}$ system at two different temperatures. (3(c)) Analogous results for the $(AB)_{6}$ systems for four different temperatures. In panels (3(b)) and (3(c)) characteristic snapshots of the final configurations are also shown as insets.
Figure 4: The phase diagram of the $(AB)_{6}$ system, displaying the density of the fluid, gas and liquid phases as extracted from direct-coexistence simulations. Note that the density of the gas at the lowest temperature ($10^{\circ}$ C) is approximated to zero, since no strand detached from the interfaces throughout the simulation. Note that the names ”gas phase” and ”liquid phase” are the ”implicit solvent” analog of polymer poor and polymer rich (network) phases.
We run very long simulations (up to $\approx 3\times 10^{10}$ simulation steps, corresponding to 2-3 months of simulation time on a NVIDIA A100 GPU), using the fraction of inter- and intra-molecular bonds as a proxy for equilibration, which we assume to be attained once these fractions reach a plateau value. We then compute density profiles along the elongated box, and classify each system as phase separated or homogeneous depending on whether the equilibrated profiles show the presence of interfaces or not.
@@
From the value of the density in the dense and in the dilute phases at equilibrium it is possible to estimate the two coexisting densities. To do so in the phase-separated samples, we classify DNA strands as belonging to the interface, gas or liquid phase, making it possible to draw a phase diagram of the system, shown in Figure 4. As the temperature increases, the density of the coexisting phases approach each other, and from the absence of interfaces at $T=17.5^{\circ}$ C we can predict the critical temperature of the $(AB)_{6}$ system to lie between $15^{\circ}$ C and $17.5^{\circ}$ C. By evaluating the average between the two coexisting densities at $T=15^{\circ}$ C we can also provide a rough estimate for the critical density, $\rho_{c}\sigma^{3}=0.06$, corresponding to a strand concentration of $\approx 1$ mM or a weight density of $\approx 53$ mg/ml.
@@
As shown in Fig. 4, for $T=15^{\circ}$ C the fraction of DNA strands in the gas is not negligible. This makes it possible to compare the properties of the single strands that make up the two phases for this particular state point. In particular, we investigate the difference in binding properties of the two phases. To this end, we show in Figure 5 the histograms of the inter-strand, intra-strand and total sticky sequences involved in bonds. The zig-zag pattern in the intra-strand histogram originates from the fact that each intra-strand bond involves two sticky sequences. As expected, the low density phase is rich in intra-molecular bonds and poor in inter-molecular bonds, while the reverse is true for the high density phase. At this temperature, which is the only one for which we have a sufficient number of gas strands to analyse, the two phases slightly differ in the total number of bonds ($\approx 9$ and $\approx 7.5$ average bonded stickers for the liquid and gas strands, respectively), signaling a weak energetic contribution to the difference in chemical potential. This small difference in the number of bonds will vanish at even lower temperatures, strongly suggesting that the two phases will have the same energy and thus that the separation is indeed driven by a strong entropic component stemming from the different conformations and different topology of the bonds adopted by the strands in the two phases.
@@ IV Conclusions
The results here presented not only validate previous theoretical work but also highlight the utility of DNA as a tunable platform for investigating the thermodynamics of associative networks. Our approach opens the door to the experimental realization of entropy-driven phase behavior in all-DNA systems, with potential implications for the design of responsive materials and hopefully for a deeper understanding of systems where phase separation of biomacromolecules is important [9, 26].

### v2

# Entropy-driven phase behaviour of all-DNA associative polymers

@@ I Introduction
In recent years, associative polymers have been extensively employed as model systems to understand the behavior of a diverse range of biopolymers comprising nucleic acids and proteins that can bind through both specific and non-specific multivalent interactions, which have been shown to be able to drive phase separation and compartmentalization without membranes [11, 12]. To capture these complex behaviors, associative polymers are typically modeled using coarse-grained representations, where the backbone is treated as a polymer chain and the stickers as discrete sites capable of specific pairwise interactions [1, 2]. In the case of simple models, where chains are flexible, spacers interact through excluded volume only, and each sticker can be involved in no more than one bond, there is theoretical [3], numerical [13, 14, 15], and experimental [7, 16] evidence that no thermodynamic instability is encountered upon increasing concentration. Indeed, in the dilute regime most of the bonds involve stickers belonging to the same chain (intra-molecular bonds), which therefore resembles a single-chain nanoparticle [17]. However, as concentration increases, chains start to interact more strongly with each other, forming inter-molecular connections that are entropically favoured [18, 19], and therefore progressively replace the intra-molecular bonds, thereby giving raise to a network-like structure without any phase separation occurring. However, it has been recently shown that in simple bead-spring models the different entropic contributions that control the overall bonding process can be manipulated by introducing multiple sticker types, with the constraint that only like stickers can bind to each other. Interestingly, as soon as there are at least two sticker types alternating along the backbone, an entropy-driven phase separation appears for a wide range of system parameters [20, 21], akin to other soft-matter phase transitions induced by entropy, such as those due to depletion [22, 23], combinatorial [24], and shape-driven [25, 26] interactions. The main contribution dictating the phase behaviour is due to the much larger entropic cost of closing an intra-molecular bond compared to an inter-molecular bond, effectively enhancing the number of inter-molecular contacts, and therefore the overall tendency to phase separate.
@@ II Methods
We simulate two one-component model systems made of single strand DNA in solution that we will call $(AA)_{6}$ and $(AB)_{6}$. Each DNA strand is composed by $N_{\rm nucl}=158$ nucleotides, with a backbone of thymine bases (the spacers) in which are inserted $M=12$ self-complementary regions (the stickers) as shown in Fig. 1, which also reports the full sequences of the strands used throughout this work. In the $(AA)_{6}$ systems the stickers have all the same sequence (TCGA), while in the $(AB)_{6}$ systems the stickers have alternating TCGA and GTAC sequences. The two sequences are self-complementary, so that only stickers of the same type can bind to each other.
@@
The DNA strands are modelled using oxDNA [27, 28], and investigated by running molecular dynamics simulations with the oxDNA standalone package [29, 30]. oxDNA is a coarse-grained model specifically designed to simulate the behavior of DNA at the nucleotide level. In the model each nucleotide is represented as a rigid body with effective interactions that capture essential structural, mechanical, and thermodynamic properties which makes it possible to study large-scale nucleic acid systems with high computational efficiency while maintaining physical realism. For all simulations we used the sequence-dependent version of the oxDNA2 model [27], with the salt concentration set to $S=0.5$ M and the integration step to $\delta t=0.002$ in reduced oxDNA units, corresponding to $\approx 6$ fs [31]. In the following, we use the oxDNA length unit $\sigma=0.8518$ nm to rescale lengths and related quantities (e.g. density). In particular, density is expressed as the dimensionless nucleotide number density $\rho\sigma^{3}=N_{\rm nucl}N_{s}\sigma^{3}/V$, where $N_{s}$ is the number of strands and $V$ is the volume of the simulation box. We note for convenience that $\rho$ can be converted to a strand molar concentration $C_{\rm mol/l}=\rho\sigma^{3}\cdot 2.69/N_{\rm nucl}$ or to a mass concentration $C_{\rm mg/ml}=C_{\rm mol/l}\cdot M_{W}$, with the molecular weight ($M_{W}$) being $48291.6~\text{g/mol}$ for both strand types.
@@
During all simulations we keep track of pairs of nucleotides sharing hydrogen bonds (HBs), where two nucleotides are considered bonded if their hydrogen bond interaction energy is lower than 0.1-0.1 in oxDNA internal units, classifying the bond as intra- or inter-molecular if the two nucleotides belong to the same strand or to different strands, respectively. We then consider two stickers as bonded if their nucleotides are involved in at least two bonds. We thus obtain the total number of stickers involved in intra- and inter-molecular bonds, $N_{\rm intra}$ and $N_{\rm inter}$, and define the fraction of stickers involved in intra- and inter-molecular as $f_{\rm intra}=N_{\rm intra}/N_{\rm max}$ and $f_{\rm inter}=N_{\rm inter}/N_{\rm max}$, where $N_{\rm max}=MN_{s}$ is the total number of stickers. We also define the fraction of bonded stickers as $f_{\rm tot}=(N_{\rm intra}+N_{\rm inter})/N_{\rm max}$.
@@ III Results
Figure 2: The fraction of hydrogen bonds (HBs) between nucleotides belonging to stickers involved in total, intra- or inter-molecular bonds for the (left) $(AA)_{6}$ and (right) $(AB)_{6}$ systems. From top to bottom, rows correspond to densities $\rho\sigma^{3}=0.57$, $\rho\sigma^{3}=0.31$, and $\rho\sigma^{3}=0.16$.
@@
Figure 3: The average end-to-end distance of homogeneous (AA)6 (full symbols joined by solid lines) and (AB)6 (empty symbols joined by dashed lines) systems as a function of $\rho$, for the three lowest temperatures simulated. The error bars have been estimated by dividing the standard deviation of the mean by $\sqrt{N_{s}}=\sqrt{128}$, which provides an upper bound for the error.
In order to understand how the interplay between total, inter- and intra-molecular bonds affect the conformation of the strands, we report in Figure 3 the average end-to-end distance of the low-temperature systems.
For the (AA)6 systems, the end-to-end distance $r_{\mathrm{ee}}$ displays minimal dependence on temperature and only a weak dependence on density. Specifically, $r_{\mathrm{ee}}$ increases slightly with increasing $\rho$, which we attribute to a gradual shift from intra- to inter-molecular bonding as the system becomes denser.
The behavior of the (AB)6 systems is somewhat more nuanced. While temperature again has a weak effect, $r_{\mathrm{ee}}$ decreases significantly with increasing $\rho$. This trend is likely driven by the growing total number of bonds as $\rho$ increases (see Figure 2), which leads to a more compact molecular conformation.
Notably, at the highest density considered, the end-to-end distance of the (AA)6 systems exceeds that of the (AB)6 systems. As shown in Figure 2, the total number of bonds is nearly identical for both strand types at this density. Therefore, the observed difference in $r_{\mathrm{ee}}$ can be attributed to the nature of the intra-molecular bonding: in (AB)6 systems, these bonds result in more compact configurations, indicating a higher entropic cost compared to their (AA)6 counterparts.
@@
Figure 4: (4(a)) Graphic representation of the system in a direct-coexistence simulation at the beginning of the calculation. (4(b)) Density profiles at the start of the simulation (dashed black curve) and at the end of the simulation, when the system has reached equilibrium (solid orange), for the $(AA)_{6}$ system at two different temperatures. (4(c)) Analogous results for the $(AB)_{6}$ systems for four different temperatures. In panels (4(b)) and (4(c)) characteristic snapshots of the final configurations are also shown as insets. All the direct-coexistence simulations shown here have been initialised from homogeneous starting configurations with $\rho\sigma^{3}=0.57$.
Figure 5: The phase diagram of the $(AB)_{6}$ system, displaying the density of the fluid, gas and liquid phases (i.e. state points where phase separation does not occur) as extracted from direct-coexistence simulations. Note that the density of the gas at the lowest temperature ($10^{\circ}$ C) is approximated to zero, since no strand detached from the interfaces throughout the simulation. Note that the names “gas phase“ and “liquid phase” are the ”implicit solvent” analog of polymer poor and polymer rich (network) phases.
As shown in the Appendix, the characteristic bond-bond autocorrelation time is $\approx 10^{9}$ time steps at $15^{\circ}$ C, and therefore we run simulations that spans longer time frames (up to $\approx 3\times 10^{10}$ simulation steps, corresponding to 2-3 months of simulation time on a NVIDIA A100 GPU). We use the fraction of inter- and intra-molecular bonds as a proxy for equilibration, which we assume to be attained once these fractions reach a plateau value. We then compute density profiles along the elongated box, and classify each system as phase separated or homogeneous depending on whether the equilibrated profiles show the presence of interfaces or not.
@@
From the value of the density in the dense and in the dilute phases at equilibrium it is possible to estimate the two coexisting densities. To do so in the phase-separated samples, we classify DNA strands as belonging to the interface, gas or liquid phase, making it possible to draw a phase diagram of the system, shown in Figure 5, which also shows studied state points where the interfaces melted and the density became homogeneous during the course of the direct-coexistence simulation (green points labelled as “fluid”). As the temperature increases, the density of the coexisting phases approach each other, and from the absence of interfaces at $T=17.5^{\circ}$ C we can predict the critical temperature of the $(AB)_{6}$ system to lie between $15^{\circ}$ C and $17.5^{\circ}$ C. By evaluating the average between the two coexisting densities at $T=15^{\circ}$ C we can also provide a rough estimate for the critical density, $\rho_{c}\sigma^{3}=0.06$, corresponding to a strand concentration of $\approx 1$ mM or a weight density of $\approx 53$ mg/ml. This value is of the same order of magnitude of the overlap concentration $\rho^{*}\sigma^{3}=N_{\rm nucl}/R_{g}^{3}\approx 0.18$ (i.e. $\approx 3$ mM strand concentration), as estimated by using the data of the coexisting liquid phase at $T=15^{\circ}$ C (see below for a more thorough discussion), for which we find $R_{g}\approx 9.5\,\sigma$. This is in line with expectations, since in associative polymer systems chain-chain interactions becomes important around the overlap concentration [3].
@@
As shown in Fig. 5, for $T=15^{\circ}$ C the fraction of DNA strands in the gas is not negligible. This makes it possible to compare the properties of the single strands that make up the two phases for this particular state point. In particular, we investigate the difference in binding properties of the two phases. While the current number of gas-phase strands provides only preliminary statistical insights, we aim to enhance these results in future work by performing targeted simulations of the two phases.
Figure 6 shows the histograms of the inter-strand, intra-strand and total sticky sequences involved in bonds. The zig-zag pattern in the intra-strand histogram originates from the fact that each intra-strand bond involves two sticky sequences. As expected, the low density phase is rich in intra-molecular bonds and poor in inter-molecular bonds, while the reverse is true for the high density phase. At this temperature, which is the only one for which we have a sufficient number of gas strands to analyse, the two phases slightly differ in the total number of bonds ($\approx 9$ and $\approx 7.5$ average bonded stickers for the liquid and gas strands, respectively), signaling a weak energetic contribution to the difference in chemical potential. Since, as shown in Fig. 2, the fraction of formed bonds increases upon decreasing temperature regardless of the density, we expect that the observed small difference in the number of bonds will vanish at temperatures lower than those investigated here, strongly suggesting that the two phases will have the same energy and thus that the separation is indeed driven (especially at low temperature) by a strong entropic component stemming from the different conformations and different topology of the bonds adopted by the strands in the two phases.
@@ IV Conclusions
The results here presented not only validate previous work but also highlight the utility of DNA as a tunable platform for investigating the thermodynamics of associative networks. Our approach opens the door to the experimental realization of entropy-driven phase behavior in all-DNA systems, with potential implications for the design of responsive materials and hopefully for a deeper understanding of systems where phase separation of biomacromolecules is considered to be important [9, 32].
Acknowledgements
We acknowledge support by ICSC – Centro Nazionale di Ricerca in High Performance Computing, Big Data and Quantum Computing, funded by European Union - NextGenerationEU, and CINECA-ISCRA for HPC resources. L.R acknowledges support from MUR-PRIN Grants No. 20225NPY8P and P2022JZEJR, funded by European Union - NextGenerationEU, Missione 4 Componente 2 - CUP B53D23028570001.
Data availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Appendix A The bond-bond autocorrelation function
Figure 8: The bond-bond autocorrelation function for the (AA)6 and the (AB)6 systems in an homogeneous state point (for densities listed in the label), and for the liquid component of the (AB)6 in the direct-coexistence simulation. All data refer to $T=15^{\circ}$ C. The crossing between the data and the horizontal orange dashed line (drawn at a value $1/e$) is the conventional estimate for the characteristic decorrelation time.
Figure 8 shows the bond-bond autocorrelation function $B(t)$, which quantifies the fraction of bonds that were present at time $t_{0}$ and are still present at time $t_{0}+t$, averaged over $t_{0}$. The figure shows that at $T=15^{\circ}$ C, for which we can compare results from the homogeneous and phase-separated samples, the characteristic time scale associated to bonding, defined as the time at which $B(t)=1/e$, is of the order of $\approx 10^{9}$ time steps.